Circular No. 7827

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 WF_{2} (LONEOS) An apparently asteroidal object discovered by LONEOS (discovery ob-

servation below) and designated 2001 WF_{2} (cf. MPEC 2001-W42) has been found to have a well-defined 45′′ tail in p.a. 320^{o} on CCD images obtained on Feb. 13.5 UT by T. B. Spahr with the 1.2-m reflector at Mount Hopkins. Following notification by Spahr, C. W. Hergenrother also found a 27′′ tail in p.a. 320^{o} and a stellar central condensation on a 1500-s co-added R-band image taken with the Catalina 1.54-m reflector.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Nov. 17.27290  3^{h}17^{m}53.^{s}91  −5^{o}02′05′′.4  18.9

Additional astrometry, the following orbital elements, and an ephemeris appear on MPEC 2002-C94:

Epoch = 2002 Feb. 15.0 TT

T = 2002 Jan. 29.84804 TT  ω = 51.^{o}35236
e = 0.6667243  Ω = 75.13262 }2000.0
q = 0.9763511 AU  i = 16.92268

a = 2.9295595 AU n^{o} = 0.19656246 P = 5.014 years

S/2002 (3749) 1 W. J. Merline, Southwest Research Institute; L. M. Close and N.

Siegler, University of Arizona; C. Dumas, Jet Propulsion Laboratory; C. Chapman, Southwest Research Institute; F. Rigaut, Gemini Observatory; F. Menard, Observatoire de Grenoble; W. M. Owen, Jet Propulsion Laboratory; and D. C. Slater, Southwest Research Institute, report the discovery on Feb. 8.4 UT of a satellite of (3749) Balam (V = 15.5; estimated diameter 7 km) from J-, H-, and K′-band direct imaging with the 8.1-m Gemini North Telescope (+ Hokupa’a adaptive optics system) on Mauna Kea: “The satellite shows no detectable motion over a 3.1-hr baseline, consistent with the expected period (of order 80 days) for this large orbit. A search for known background/foreground small bodies with similar position and velocity reveals no candidates brighter than mag 21. The brightness ratio in band H is ∼ 3.3 mag, giving an estimated diameter of the satellite of ∼ 1.5 km. On Feb. 8.4311, the secondary was at 0′′.47 separation with position angle 115^{o}.”

2002 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green